Explain why the equation x(x+4) =12 cannot be solved by solving the two equations x=12 and x+4=12.
1
Expert's answer
2019-02-28T09:57:07-0500

If x = 12, x+4 = 12, then x(x+4) = 144, but it is given that the initial equation is x(x+4) = 12. We got a contradiction, hence assumption 'x=12, x+4=12' was wrong. Thus, the equation x(x+4) =12 cannot be solved by solving the two equations x=12 and x+4=12.